If 3x and (2x-5) are the measure for complementary angles, what is the measure of each angle?
Answer by stanbon(48529)

If 3x and (2x-5) are the measure for complementary angles, what is the measure of each angle?
---
The sum of complementary angles is 90 degrees
Equation:
3x + 2x-5 = 90
5x = 95
x = 19
-----
3x = 57 degrees (one of the angles)
2x-5 = 2*19-5 = 38-5 = 33 degrees (the other angle)
=================================================
Cheers,
Stan H.